The following table summarizes the collected data from 25 quadrats in each area

Species

Total number in area 1

Rel. abundance in area

Total number in area 2

Relative abundance in area 2

 

1

Artemisia herba-alba

44

0.076125

35

0.093582888

2

Ballota undulata

10

0.017301

0

0

3

Salvia dominica

25

0.020761

1

0.002673797

4

Carlina hispanica

12

0.020761

2

0.005347594

5

Ononis natrix

14

0.024221

0

0

6

Poa bulbosa

350

0.041522

.

#VALUE!

7

Stipa barbata

24

0.041522

20

0.053475936

8

Salsola vermiculata

48

0.083045

250

0.668449198

9

Astragauls spinosus

2

0.00346

6

0.016042781

10

Astragalus callichrous

22

0.038062

35

0.093582888

11

Silene colorata

26

0.044983

12

0.032085561

12

Anabasis articulata

1

0.00173

13

0.034759358

Total

578

100%

374

100%

0.413495

1.898395722

 

 

 

Q. 2) Calculate the species richness, evenness and the diversity (use Shannon index) in each area (Community ecology). Compare and contrast both areas. 

1
Expert's answer
2021-05-10T07:09:02-0400

Species richness (S) = number of species of taxonomy of interest = 12 species


Area 1

Proportion of each species "(P_i)"


P of artemisia = 44/578 = 0.076125

ln P = -2.5754

P × ln(P) =-0.196


P of ballota = 10/578 = 0.017301

ln P = -4.057

P × ln(P) = -0.0702


P of Salvia = 25/578 = 0.04325

ln P = -3.140

P × ln(P) = -0.1358


P of carlina = 12/578 = 0.0207

ln P = -3.878

P × ln(P) = -0.0804


P of ononis = 14/578 = 0.024221

ln P = -3.7205

P × ln(P) = -0.0901


P of poa = 350/578 = 0.6055

ln P = -0.5017

P × ln(P) = -0.3038


P of stipa = 24/578 = 0.041522

ln P = -3.1815

P × ln(P) = -0.132

l

P of salsola = 48/578 = 0.083045

ln P = -2.488

P × ln(P) = -0.2066


P of astragamus = 2/578 = 0.00346

ln P = -5.667

P × ln(P) = -0.0196


P of astragalus = 22/578 = 0.44983

ln P = -3.2685

P × ln(P) = -0.1244


P of silene = 26/578 = 0.044983

ln P = -3.102

P × ln(P) = -0.1395


P of anabasis = 1/578 = 0.00172

ln P = -6.360

P × ln(P) = -0.011003



Shannon’s diversity index (H) = "- \Sigma [P_i×\ln P_{i}]= -0.196 -0.072 -0.1358 -0.0804-0.0901 - 0.3038 -0.132 -0.2066 -0.0196 -0.1244 -0.1395-0.011003 = -1.511203"


Species Evenness "=\dfrac{1.511203}{12}= 0.1259"